OpenStudy (anonymous):

Please help me with this arithmetic series question quick? Thanks! Find the sum: 28 is on top of the E, I=1 is on the bottom of the E, and (8i - 13) is on the right side of the E. 2884 205 3024 211

ganeshie8 (ganeshie8):

\(\large \sum \limits_{i=1}^{28} (8i-13)\)

ganeshie8 (ganeshie8):

like that ?

OpenStudy (anonymous):

Yes :)

ganeshie8 (ganeshie8):

remember the "arithmetic sum " formula ?

ganeshie8 (ganeshie8):

Sum of n terms in arithmetic sequence, whose first term is \(a\) and last term \(l\) : \(S_n = \frac{n}{2}(a+l)\)

ganeshie8 (ganeshie8):

for the given series, clearly \(n=28\)

OpenStudy (anonymous):

Okay. S I know that n would = 28 but then I'm lost after that

ganeshie8 (ganeshie8):

if u can find \(a\) and \(l\), u can simply use that sum formula

OpenStudy (anonymous):

So. A1 would = -5

OpenStudy (anonymous):

@Eckwright576 Also remember the generalized formula for sum of an arithmetic series: S=(n/2)*[2a+(n-1)*d] where, d is the common difference, a is the first term and n is the no of term.

ganeshie8 (ganeshie8):

yes, first term is \(-5\)

ganeshie8 (ganeshie8):

how about last term ?

ganeshie8 (ganeshie8):

here, last term is 28th term right ? u need to find \(a_{28}\)

ganeshie8 (ganeshie8):

plugin \(i = 28 \) in \(8i - 13\)

ganeshie8 (ganeshie8):

that gives u 28th term

OpenStudy (anonymous):

-140?

ganeshie8 (ganeshie8):

try again

ganeshie8 (ganeshie8):

8(28) - 13

ganeshie8 (ganeshie8):

simplify

OpenStudy (anonymous):

211

ganeshie8 (ganeshie8):

yes ! so u have \(a = -5\) \(l = 211\)

ganeshie8 (ganeshie8):

you're ready to use the sum formula ?

OpenStudy (anonymous):

Yeah. I think

ganeshie8 (ganeshie8):

good, plug them and see wat u get for sum

OpenStudy (anonymous):

I got 2575. But that's none of my answers?

ganeshie8 (ganeshie8):

try again, u must have done a calculation mistake

ganeshie8 (ganeshie8):

\(S_n = \frac{n}{2}(a+l)\) \(S_{28} = \frac{28}{2}(-5+211)\) \(~~~~= 14(206)\) \(~~~~= ?\)

OpenStudy (anonymous):

I figured out my mistake! :) the answer is. 2884

ganeshie8 (ganeshie8):

\(\large \color{red}{\checkmark}\)